Show that the set {1 βˆ’ π‘₯, 3 βˆ’ π‘₯ 2 , π‘₯} spans 𝑃2


1
Expert's answer
2022-03-05T07:27:16-0500

Solution


For an arbitrary polynomial of degree 2,


"p(x)=ax^2+bx+c"


To prove that S is a spanning set, we need to show that "p(x)" can be written as linear combination of the elements of "S=\left \{ 1-x, 3-x^2,x \right \}".


Then, for the three constants, "r", "s" and "t" , we can write


"p\left( x \right) = r\left( {1 - x} \right) + s\left( {3 - {x^2}} \right) + t\left( x \right)\\
a{x^2} + bx + c = r - rx + 3s - s{x^2} + tx\\
a{x^2} + bx + c = - s{x^2} + \left( { - r + t} \right)x + \left( {r + 3s} \right)"


This gives,


"a = - s \Rightarrow a= 0 \cdot r - 1 \cdot s + 0 \cdot t\\

b = - r + t \Rightarrow b= - 1 \cdot r + 0 \cdot s + 1 \cdot t\\

c = r + 3s \Rightarrow c = 1 \cdot r + 3 \cdot s + 0 \cdot t\\"


To check that a solution exists, we set up the augmented matrix and row reduce it,Β 


"\begin{bmatrix}
 0 & -1 & 0 & r\\
 -1 & 0 & 1 & s\\
1 & 3 & 0 & t 
\end{bmatrix}"


which is row reduced to


"\begin{bmatrix}
 1 & 0 & 0 & 3r+t\\
 0 & 1 & 0 & -r\\
0 & 0 & 1 & 3r+s+t 
\end{bmatrix}"


Hence we have


"a=3r+t\\
b=-r\\
c=3r+s+t"


Therefore, we conclude that Therefore, S is a spanning set for P2